The reaction:

2 NO2(g) ↔ N2O4(g)

has an equilibrium constant, Kc, of 170 at 298K. Analysis of this system at 298K reveals that 4.57E-1 mol of NO2 and 6.95E-2 mol of N2O4 are present in a 31.0-L flask. Determine the reaction quotient, Q, for this mixture. Enter your answer in scientific notation.

Which direction must the reaction in the previous question proceed to reach equilibrium. Note that you have only two tries for this question!

A. Right
B. Left
Answer #1


1)
[NO2] = mol of NO2 / volume in L
= 4.75*10^-1 mol / 31.0 L
= 1.53*10^-2 M

[N2O4] = mol of N2O4 / volume in L
= 6.95*10^-2 mol / 31.0 L
= 2.24*10^-3 M

Now use:
Qc = [N2O4]/[NO2]^2
= (2.24*10^-3) / (1.53*10^-2)^2
= 9.57

Answer: 9.57

2)
Qc = 9.57
Kc = 170
Qc is less than Kc.
So, reaction proceed in forward direction that is to the right
Answer: Right
Answer #2

Step 1: Calculate the Reaction Quotient (Q)

The reaction quotient Q is calculated similarly to the equilibrium constant Kc, but uses the current concentrations instead of equilibrium concentrations.

Given Reaction:

2NO2(g)N2O4(g)

Formula for Q:

Q=[N2O4][NO2]2

Given Data:

  • Moles of NO2=4.57×101mol

  • Moles of N2O4=6.95×102mol

  • Volume V=31.0L

Calculate Concentrations:

  1. [NO2]=4.57×101mol31.0L=1.474×102M

  2. [N2O4]=6.95×102mol31.0L=2.242×103M

Compute Q:

Q=2.242×103(1.474×102)2=2.242×1032.173×104=10.32

Scientific Notation:

Q=1.03×101


Step 2: Compare Q to Kc to Determine Reaction Direction

  • Given Kc=170 (equilibrium constant).

  • Calculated Q=10.32.

Key Rule:

  • If Q<Kc: Reaction proceeds right (toward products, N2O4).

  • If Q>Kc: Reaction proceeds left (toward reactants, NO2).

Here, Q(10.32)<Kc(170), so the reaction shifts:
A. Right (to form more N2O4 until Q=Kc).


Answers:

  1. Reaction Quotient (Q): 1.03×101

  2. Direction to Reach Equilibrium: A. Right


answered by: anonymous